# Foundation Models in Smart Agriculture: Basics, Opportunities, and Challenges

### 21 Dec 2023 v3

@@ 4.3 Distribution shift
To combat the challenges posed by distribution shifts when training and deploying FMs, researchers have been continuously investigating these issues (Chawla et al., 2021, Cvejoski et al., 2022, Bommasani et al., 2021). For instance, Cvejoski et al. (2022) demonstrated that Large Language Models (LLMs) may experience performance drops of approximately 88% at best when predicting the popularity of future posts from subreddits whose topic distribution evolves over time. To mitigate these distribution shift issues, they introduced neural variational dynamic topic models and attention mechanisms. These techniques infer temporal language model representations for regression tasks, resulting in only about 40% performance drops in the worst-case scenarios and a mere 2% in the best cases. These techniques have been shown to significantly reduce performance drops caused by distribution shifts, thus demonstrating promising prospects for the application of FMs in the diverse and dynamic field of agriculture.
@@ 4.4 Other challenges in real-world deployment
Deploying FMs for agricultural applications presents a unique set of challenges, particularly regarding inference speed and model size. Firstly, the inherent complexity and size of these models can pose substantial computational demands. Large-scale FMs often require significant processing power and memory for both training and inference, which may not always be readily available in on-field agricultural settings (Bommasani et al., 2021). This can limit the deployment of these models in real-time applications, such as monitoring crops or automating irrigation systems. Secondly, inference speed is crucial for many agricultural tasks that require immediate action based on the model’s outputs (Bommasani et al., 2021). For instance, in precision agriculture, time-sensitive tasks like pest detection or predicting weather impacts on crop health demand quick and accurate responses. However, due to the size and complexity of foundation models, achieving high-speed inference can be a challenge.
@@
To overcome these hurdles, future research and development efforts need to focus on model optimization techniques (Zhong et al., 2023), such as model compression (e.g., pruning and quantization (Cheng et al., 2017, Choudhary et al., 2020)) and efficient network design (Wan, 2022), which could potentially reduce the model size without compromising performance. Actually, there are already some attempts for deploying FMs. For instance, a fully quantized vision transformer, FQ-ViT (Lin et al., 2022), was proposed to simplify inference by using only 4-bit quantization and the BitShift operator with only nearly lossless accuracy degradation ($\sim 1\%$). SqueezeLLM (Kim et al., 2023) showed that 3-bit quantization significantly reduces the perplexity gap from the FP16 baseline by up to 2.1 times and achieved up to 2.3 times speedup compared to the baseline models.
@@
Additionally, the utilization of edge computing solutions (Chen and Ran, 2019) could expedite the inference process by enabling data processing closer to its source, consequently enhancing the speed and efficiency of these models in agricultural applications. Recently, MLC LLM (Girshick et al., 2023) offers a universal solution, allowing any language model to be deployed natively across a diverse range of hardware backends and native applications. This solution also provides a productive framework for individuals to further optimize model performance for their specific use cases. On the other hand, MobileSAM (Zhang et al., 2023b) made SAM (Kirillov et al., 2023) mobile-friendly by replacing the heavyweight image encoder with a lightweight one. This lightweight SAM is more than 60 times smaller yet performs on par with the original SAM (Kirillov et al., 2023).
5 Summary
@@
In conclusion, this study has investigated the potential of foundation models in the field of smart agriculture. These models hold the promise of transforming the agricultural sector by reducing reliance on extensive labeled datasets and enhancing generalization and effectiveness. By categorizing and reviewing various types of foundation models (Section 2), we have revealed their applicability in smart crop management, smart plant breeding, smart livestock farming , precision aquacuture farming , and agricultural robotics (Section 3). However, their utilization in smart agriculture is still in its early stages, and challenges and risks such as model training, validation, and deployment need careful consideration (Section 4). Further exploration and utilization of foundation models have the potential to enhance productivity, sustainability, and decision-making in agriculture. This study aims to inspire and guide researchers to unlock the untapped potential of foundation models in smart agriculture. We hope that this manuscript can serve as a stepping stone toward developing smart agriculture with next-level intelligence and more sophisticated capabilities.
@@
Jiajia Li: Conceptualization, Investigation, Software, Writing – original draft; Mingle Xu: Conceptualization, Investigation, Writing – original draft; Lirong Xiang: Investi-
gation, Writing - review & editing; Dong Chen: Conceptualization, Investigation, Software, Writing – original draft; Weichao Zhuang: Writing - review & editing; Xunyuan Yin: Writing - review & editing; Zhaojian Li: Supervision, Writing - review & editing.

### 17 Mar 2024 v4

@@ 4.3 Distribution shift
To combat the challenges posed by distribution shifts when training and deploying FMs, researchers have been continuously investigating these issues (Chawla et al., 2021, Cvejoski et al., 2022, Bommasani et al., 2021). For instance, Cvejoski et al. (2022) demonstrated that LLMs may experience performance degradation of up to approximately 88% when predicting the popularity of future posts from subreddits whose topic distribution evolves over time. To mitigate these distribution shift issues, they introduced neural variational dynamic topic models and attention mechanisms. These techniques infer temporal language model representations for regression tasks, resulting in only about 40% performance drops in the worst-case scenarios and a mere 2% in the best cases. Furthermore, Wang et al. (2023) demonstrated that the integration of techniques such as multi-task learning, continual learning, and distillation with VFMs like SAM and CLIP can significantly alleviate distribution shift issues. These techniques have been shown to significantly mitigate performance degradation caused by distribution shifts, thus demonstrating promising prospects for the application of FMs in the diverse and dynamic field of agriculture.
@@ 4.4 Other challenges in real-world deployment
Deploying FMs for agricultural applications presents a unique set of challenges, particularly regarding inference speed and model size. Firstly, the inherent complexity and size of these models can pose substantial computational demands. Large-scale FMs often require significant processing power and memory for both training and inference, which may not always be readily available in on-field agricultural settings (Bommasani et al., 2021). This can limit the deployment of these models in real-time applications, such as monitoring crops or automating irrigation systems. Secondly, inference speed is crucial for many agricultural tasks that require immediate action based on the model’s outputs (Bommasani et al., 2021). For instance, in precision agriculture, time-sensitive tasks like pest detection or predicting weather impacts on crop health demand quick and accurate responses. However, due to the size and complexity of FMs, achieving high-speed inference can be a challenge.
@@
To overcome these hurdles, future research and development efforts need to focus on model optimization techniques (Zhong et al., 2023, Zhu et al., 2023), such as model compression (e.g., pruning and quantization (Cheng et al., 2017, Choudhary et al., 2020)) and efficient network design (Wan, 2022), which could potentially reduce the model size without compromising performance. Actually, there are already some attempts to deploy FMs. For instance, FQ-ViT (Lin et al., 2022) introduces a fully quantized vision transformer that simplifies inference by utilizing only 1/4 of the memory with minimal accuracy degradation (approximately 1%). Additionally, SqueezeLLM (Kim et al., 2023) demonstrates that utilizing only 3/16 of the parameters significantly reduces the perplexity gap by up to 2.1 times and achieves up to 2.3 times speedup compared to baseline models.
@@
Additionally, the utilization of edge computing solutions could expedite the inference process by enabling data processing closer to its source, consequently enhancing the speed and efficiency of FMs in agricultural applications (Chen and Ran, 2019). For example, in agricultural applications, edge devices like drones or field sensors can analyze data locally, enabling real-time decision-making without relying on distant servers. Recently, MLC LLM (Girshick et al., 2023) offers a universal solution, allowing any language model to be deployed directly across a diverse range of hardware backends and native applications. This solution also provides a productive framework for individuals to further optimize model performance for their specific use cases. On the other hand, MobileSAM (Zhang et al., 2023b) and MobileSAMv2 (Zhang et al., 2023c) make SAM (Kirillov et al., 2023) mobile-friendly by replacing the heavyweight image encoder with a lightweight one. This lightweight SAM is more than 60 times smaller yet performs on par with the original SAM (Kirillov et al., 2023).
4.5 Building capacity for FM adoption in agriculture
Effective utilization of FMs in agriculture requires specialized skills and multi-discipline collaborations (Mai et al., 2023). However, there are several challenges in the skill transfer of applying FMs in agricultural domains, including domain adaptation techniques due to agricultural researchers’ potential lack of familiarity with FM applications, understanding unique characteristics of agricultural data such as high variability, and interpreting the complex nature of FMs.
Addressing the skill gap and providing targeted training is crucial for empowering agricultural researchers and users to leverage FMs optimally. Investing in training and capacity-building initiatives, such as workshops, webinars, specialized training programs focusing on FM utilization in agriculture, and educational resources such as online tutorials and case studies, is essential for enabling effective FM utilization in agriculture, driving innovation, and addressing agricultural challenges.
5 Conclusion
@@
In conclusion, this study has investigated the potential of FMs in the field of smart agriculture. These models hold the promise of transforming the agricultural sector by reducing reliance on extensive labeled datasets and enhancing generalization and effectiveness. By categorizing and reviewing various types of LLMs and FMs (Section 2), their applicability in smart crop management, smart plant breeding, smart livestock farming, precision aquaculture farming, and agricultural robotics are revealed (Section 3). However, it is crucial to address challenges and risks such as data collection, model training, distribution shift, real-world deployment, and research gaps before real-world implementation (Section 4). Further exploration and utilization of LLMs and FMs have the potential to enhance productivity, sustainability, and decision-making in agriculture.
In the era of LLMs and FMs, all agricultural fields successfully explored using traditional ML/DL approaches, as discussed in Section 3 and beyond, are ready for exploration by employing FMs. Integrating LLMs and FMs into the existing framework is also a promising direction to improve efficiency and effectiveness, as evident in the studies by Stella et al. (2023), Tzachor et al. (2023). These integrations can leverage the strengths of both frameworks, potentially leading to more robust and adaptable agricultural systems. Furthermore, the development of large-scale agricultural datasets would be an approaching direction to accelerate the development and application of LLMs and FMs in the agriculture field (Lu and Young, 2020). Such datasets can facilitate more comprehensive training of models, enabling them to capture the nuances and complexities of agricultural processes more accurately. Additionally, there is a growing need for interdisciplinary collaboration between experts in agriculture, computer science, and data science. By fostering collaboration across these domains, innovative solutions can be developed to address the unique challenges and opportunities in agriculture using LLMs and FMs.
Ultimately, this study aims to inspire and guide researchers to unlock the untapped potential of LLMs and FMs in smart agriculture and serves as a stepping stone toward developing smart agriculture with next-level intelligence and more sophisticated capabilities.
@@
Jiajia Li: Conceptualization, Investigation, Methodology, Software, Writing – original draft; Mingle Xu: Formal Analysis, Writing - review & editing; Lirong Xiang: Formal Analysis, Writing - review & editing; Dong Chen: Conceptualization, Investigation, Writing – original draft; Weichao Zhuang: Writing - review & editing; Xunyuan Yin: Writing - review & editing; Zhaojian Li: Supervision, Resources, Writing - review & editing.